# Black Hole Entropy, Quantum Corrections and EFT Transitions

### v1

@@ 3.4.2 Alternative computation of the Schwinger determinant
Subsequently, we can add to the integration contour the semi-circle at infinity in the upper half plane since it does not contribute to the complex integral. [^19] This is true, in general, only if $\text{Re}\,\alpha\neq 0$. See Section 4.3.1 for details on this point. Finally, by gluing the integrals in (3.46) to avoid the pole at the origin, we construct a closed path in the complex $s$-plane (see Figure 5), which enables us to rewrite (3.26) as follows [^20] We emphasize that a similar approach to obtaining the full contribution of a given D0-D2 bound state to the generalized prepotential was recently proposed in Hattab:2024ewk .
Figure 5: Integral contour in the complex $s$-plane employed to evaluate the integral (3.47). The singularities located along the imaginary axis are associated to non-perturbative pair supergraviton production, whereas the real poles correspond to the perturbative piece (3.28).
@@
which can be finally evaluated upon using the residue theorem. Interestingly, there are two kinds of poles that contribute to the integral (3.47). On the one hand, those occurring along the real axis $s=k\in\mathbb{Z}$ provide for the perturbative piece already computed in (3.28). On the other hand, the poles located at $s=\frac{2\pi in}{\alpha}$ account for non-perturbative pair production. Hence, as a final result one obtains
@@
Let us take the opportunity here to stress that the fact that both prescriptions to compute the perturbative and non-perturbative quantum contributions to the generalized prepotential due to the D0-brane tower agree rests, at the end of the day, on us being able to deform the contours for positive/negatively charged states, as well as to add the arc at infinity in Figure 5 without any additional cost. Importantly, though, this may not always be necessarily the case, which ultimately depends on the complex phase that the expansion parameter $\alpha$ exhibits (cf. footnote 19). We will elaborate further on this topic later on in Section 4.3.1.
@@ 4 The Fate of Other BPS Black Hole Systems
To accomplish this, we analyze in Section 4.1 another BPS system involving D2- and D6-brane charge. The reason for selecting this family of solutions will become clear along the way. Therefore, following the same strategy as in the previous chapter, we first describe these black holes from the perspective of the two-derivative supergravity theory. Subsequently, in Section 4.2, we repeat the analysis taking now into account the effect of the higher-derivative F-terms introduced around eq. (2.5). A key difference between this configuration and the D0-D2-D4 black hole system is that now the expansion parameter controlling the quantum deformations of the theory is purely imaginary. However, as we argue in Section 4.2.2, by appropriately choosing the gauge charges one is able to probe the pathological regime $|\alpha|=\mathcal{O}(1)$ here as well. Nevertheless, in Section 4.3, we show that it is still possible to include highly non-local effects due to the tower of D0-branes which allow us to resum their Schwinger contribution in an exact (i.e., non asymptotic) way. Crucially, we also find that non-perturbative effects are absent in this class of backgrounds, and hence do not modify the attractor solutions nor the entropy. This nicely matches with the predictions made in previous sections.
@@
Before proceeding with our discussion, let us briefly summarize here our findings. First of all, we observe that, as soon as we turn on the D6-brane charge in the system, the solution can no longer explore the genuine 5d regime, which we recall corresponds to the $|\alpha|\to\infty$ limit. This can be readily checked from the attractor mechanism, even at two-derivative level. The latter imposes an upper bound on the stabilized value for the (absolute value of) expansion parameter $\alpha$, whenever $p^{0}\neq 0$. Alternatively, from the M-theory perspective, one can argue that having non-trivial D6-brane charge is equivalent to introducing a background Taub-NUT geometry whose center is superimposed with the black hole singularity (see Appendix B for details). This implies, among other things, that at the attractor locus at most ${r_{h}\sim r_{5}}$ can occur, thus preventing us from performing a clean matching with the putative 5d index, similarly to what we did in Section 3.3.2. Still, by analyzing a representative example introduced in Section 4.1, we confirm that an EFT transition must happen when $|\alpha|$ becomes of order one, which is signaled by an apparent singular behavior exhibited by the asymptotic series of corrections to, e.g., the entropy. This problem can be cured by resorting to the uplifted 5d theory, thereby including highly non-local effects involving the extra circular direction. In any event, very remarkably, we find that non-perturbative contributions in the Scwhinger integral also decouple from this kind of solutions, as it was the case in the simpler D0-D2-D4 system. Finally, in Section 4.3.1 we point out that a simple alternative approach to compute the non-perturbative pair production effects based on Cauchy (cf. Section 3.4.2) seems difficult to apply herein, thus requiring from a special treatment.
@@ 4.1.1 The two-derivative analysis
The family of black hole solutions with $p^{0}=0$ introduced in Section 3.1 is rather special, since they exhibit an explicit attractor for any combination of the remaining quantized charges (see, however, footnote 11). On the other hand, in the most general situation with non-trivial $p^{0},q_{0}$ and arbitrary $(q_{a},p^{b})$-charges, the system is characterized instead by a set of algebraic quadratic equations which may or may not have a real solution, even at the classical level of approximation Shmakova:1996nz . Therefore, in order to provide yet another instance where quantum corrections to the entropy $\mathcal{S}_{\rm BH}$ can be determined —and subsequently studied, we consider in what follows the restricted case of 4d black holes with $p^{a}=0$, i.e., with no D4-brane charge. The reason for this choice is two-fold. First, because such solutions are shown to exist for any relative value of the horizon and asymptotic M-theory radii Gauntlett:2002nw ; Gaiotto:2005gf , and can be moreover easily analyzed. Secondly, since they uplift to five-dimensional, spinning BPS black holes at the center of a Taub-NUT geometry (cf. Appendix B), in contrast to the D0-D2-D4 configuration, which corresponds instead to 5d black strings extended along the compactification circle.
@@
a condition that can be easily attained upon imposing the hierarchy $q_{a}\gg p^{0}$. [^22] Notice that, given a solution $\{x^{a}\}$ of (4.4), one may obtain similar ones upon rescaling $q_{a}\to k^{\beta}q_{a}$ and ${p^{0}\to k^{\gamma}p^{0}}$, which results into $x^{a}\to k^{\frac{\beta+\gamma}{2}}x^{a}$. Hence, by taking $\beta>\gamma$ and $k\gg 1$, one can easily achieve (4.10). Notice that, similarly to what happened with the D0-D2-D4 system (cf. (3.9)), at this level of approximation we do not need to specify the asymptotics of $x^{0}$, which in the present case is entirely determined by the D6-brane charge $p^{0}$. The latter turns out to control the importance of the relevant, perturbative quantum corrections, as we discuss next.
@@ 4.2.1 Including higher-derivative corrections
$|\mathscr{Z}|^{2}$ $=-\frac{4}{3}Y^{a}\left(q_{a}-\frac{1}{12p^{0}}c_{2\,a}\right)+p^{0}\,\text{Re}\,G_{0}\,,$ (4.13)
@@
$\mathcal{S}_{\rm BH}=-\frac{4}{3}\pi Y^{a}\left(q_{a}+\frac{1}{6p^{0}}c_{2\,a}\right)+\frac{\chi_{E}(X_{3})(p^{0})^{2}}{4(2\pi)^{2}}\sum_{g=0,2,3,\ldots}(-1)^{g}c^{3}_{g-1}\,|\alpha|^{2g}\,.$ (4.14)
@@ 4.2.2 The transition regime
Let us also remark that the absence of an asymptotic five-dimensional regime prevents the existence of a genuine 5d configuration in the decompactified theory that could be used to ‘glue’ with the D2-D6 black hole across the transition regime. This observation was already clear from the perspective of the classical, two-derivative theory, given that the KK monopoles sourcing the Taub-NUT charge are topological solutions that exist only in presence of compact directions, see Appendix B.1. Furthermore, we can confirm now that the previous conclusion is not modified upon including the relevant set of quantum corrections (perturbative and non-perturbative). Therefore, we cannot test our quantum-corrected entropy in the same way as in Section 3 for the D0-D2-D4 black hole, where we ultimately recovered the (regularized) entropy of the 5d black string carrying M2, M5 and KK charges by taking the $\alpha\to\infty$ limit.
@@ 4.3 Non-local and non-perturbative effects
Figure 7: Numerical plot of $\mathcal{I}(|\alpha|)$ as a function of the (purely imaginary) expansion parameter $|\alpha|$. The actual computation corresponds to the ‘renormalized’ expression displayed in (4.22). Notice that the integral is convergent and finite for every $|\alpha|\in[0,2)$ within its physical domain (see footnote 25).
[…]
@@ 4.3.1 Challenges and Obstructions in the Cauchy formulation
$\mathcal{I}^{(np)}(\alpha)\,$  $\sim\,2\pi i\alpha\sum_{n=1}^{\infty}\frac{1}{n}e^{-\frac{4n\pi^{2}}{\alpha}}\,,$ (4.24b)
@@
Figure 9: Integral contour in the complex $s$-plane employed to evaluate the one-loop integral (4.23). The non-perturbative singularities no longer lie along the imaginary axes if $\alpha$ has a non-vanishing complex phase $\theta_{\alpha}$. The real poles still correspond to the perturbative piece. In the limit of $\text{Re}\,\alpha\to 0$, all the poles therefore become real. However, for the rational values of $\alpha$ enforced by the attractor mechanism, perturbative and non-perturbative singularities do not match.
The divergences we are encountering in the case of $\text{Re}(\alpha)=0$ signal that we should not be allowed to use anymore the residue theorem when evaluating (4.23). Indeed, in its standard formulation (see, e.g., Ahlfors1966 ), one considers a finite arc as well as finitely many poles. One can then formally extend the computation to the case of infinitely many isolated poles by considering a discrete family of contours $\{\mathcal{C}_{N}\}$ which enclose $N$ such singularities, and subsequently take the limit $N\to\infty$. This process is well-defined if and only if the limit exists, i.e., provided the series of residues converges. For $\alpha$ imaginary this does not happen, as we just discussed, and thus we cannot use the residue theorem to evaluate the integral. Notice that this oscillatory behavior can be ultimately traced back to the contributions associated to the infinitesimal semi-circles surrounding each pole along the real axis. In essence, what happens is that, since both sets of singularities are dense with respect to each other, the small arcs around the ‘quasi-poles’ get an enhancement —due to the closeness to the nearest singularity— that grows as we go towards real-infinity, as per (4.26).
In addition, it is worth highlighting that, from this perspective, one can also understand the origin of the harmonic-like behavior exhibited by the series of residues (4.25). Recall that the starting point was eq. (3.26), which does require special care when $\alpha$ is purely imaginary. Whenever this happens, what one can do is to consider a shift so as to move the non-perturbative singularities infinitesimally away from the real axis, as we did in (3.46). While this deformation may seem sufficient to eliminate the divergences caused by going exactly through poles, it is not enough to fully regularize the expression. To see this, let us consider the integrand of (3.46a) with $\alpha=i|\alpha|$, with the non-perturbative poles properly shifted as explained before, namely
@@
where we substituted $\alpha=i|\alpha|$ and we defined $\tau=is$, cf. footnote 14. Note that the above expression is clearly reminiscent of the exact result derived in eq. (4.22), and it is moreover well-defined and convergent (except for the pole at $\tau=0$, which must be carefully dealt with as in the rest of this work). Furthermore, from this point of view it becomes clear that the pathologies we referred to in our previous discussion would be absent in (4.30) and hence only appear if we close the contour by adding the arc at infinity, whereby picking up the residues displayed in (3.48).
@@ 5 Conclusions and Outlook
In this work, we have revisited the role played by macroscopic quantum corrections to the black hole entropy in 4d $\mathcal{N}=2$ supersymmetric effective field theories arising from Type IIA string theory compactified on Calabi–Yau threefolds. More precisely, by considering an infinite series of higher-derivative F-terms, we examined the most relevant perturbative and non-perturbative contributions to the supersymmetric black hole index close to the large volume point, emphasizing the interplay between different limiting (dual) descriptions in four and five spacetime dimensions. In particular, we obtained an explicit and well-defined, analytic expression for the entropy which allows us to track the underlying physical systems across the transition between the two different field-theoretic descriptions. As a byproduct, this study highlights the importance of EFT transitions in understanding quantum-gravitational aspects of black hole thermodynamics.
@@
One of the main conclusions of our investigation is that the aforementioned quantum corrections to the entropy exhibit distinct behaviors as the characteristic size of the black hole approaches the Kaluza-Klein scale, where the effects of the extra dimensions become relevant. In particular, we identified a ‘transition regime’ (cf. Section 3.2.2) where an infinite number of local higher-curvature and higher-derivative operators seem to induce pathological contributions to the macroscopic black hole entropy, signaling the failure of the 4d $\mathcal{N}=2$ EFT to correctly describe such configurations. Indeed, the corrections organize into an asymptotic series with expansion parameter $\alpha$ that can be interpreted as the ratio of the inverse D0-brane mass to the black hole radius. We showed how one can regularize the series trough a resummation procedure and how the highly non-local perturbative effects induced by the infinite tower of Kaluza-Klein states (identified herein as D0 bound states) plays a crucial role in regulating certain ultra-violet divergences naively exhibited by the entropy function. This was explicitly analyzed for two BPS configurations, namely the D0-D2-D4 (Section 3) and the D2-D6 (Section 4) black holes. Furthermore, in a sense, these two systems are somewhat complementary. From a more technical point of view, the resummation procedure that was used for the former black hole can be readily extended for more general charge configurations. Crucialy, though, it fails precisely when we turn off the D0-brane charge, which is equivalent to the regime of purely imaginary $\alpha$, see discussion in Section 4.3.1. Therefore, in Section 4, we not only studied a second example but also clarified how to extend the resummation prescription employed in Section 3 to this case as well. Moreover, as discussed in Section 3.3.2, by lifting the solution to five dimensions and enforcing a parametrically small horizon radius compared to the size of the internal circle (whenever possible), one finds perfect agreement with the microscopic counting Maldacena:1997de ; Vafa:1997gr ; Harvey:1998bx and one-loop exact computations in the dual M-theory description Castro:2007sd . The transition regime can be crossed completely by taking some appropriate limit only for the D0-D2-D4 system. Indeed, the aforementioned black hole becomes a 5d black string wrapped along the extra compact direction, which grows indefinitely in the decompactification limit. Instead, the D2-D6 solutions uplifts to a five-dimensional black hole with Kaluza–Klein monopole charge, and it exists as long as the extra dimension is strictly compact. In this latter case, we can explore the transition regime; however, a topological obstruction prevents us from taking the full decompactification limit for this class of configurations. Remarkably, we also found that additional non-perturbative corrections —which are oftentimes present and in this case may be understood as a non-trivial pair production rate of Kaluza-Klein gravitons (and superpartners thereof) in the anti-self-dual graviphoton constant background close to the black hole horizon— seem to not modify the most relevant physical properties of the solutions considered on this work, in particular their associated entropy (cf. Sections 3.4 and 4.3). On the one hand, this further supports the consistency of the analysis carried out here. On the other hand, it leads us to conjecture that the same phenomenon will occur for more general BPS black holes that can be constructed in the 4d theory. A more detailed study of these issues is left for future work.
@@
On another note, one may hope to be able to obtain from this perspective further insights into the non-perturbative behavior of certain topological string theories, which are known to capture the same prepotential controlling the higher-derivative corrections to the entropy Ooguri:2004zv . In fact, we saw that in the D0-D2-D4 black hole background, namely the one associated to the unique system that is able to explore the genuine 5d regime, there should be a priori certain non-perturbative contributions to the generalized prepotential. These can be equivalently determined via a careful study of a one-loop integral in M-theory Gopakumar:1998ii ; Gopakumar:1998jq ; Pasquetti:2010bps (see also Hattab:2024ewk ; Hattab:2024ssg for recent related results). Hence, even though the latter did not ultimately affect any of the black hole observables we actually cared about in this work, they certainly exhibited interesting behaviors, particularly so along the decompactification limit. Interestingly, this regime can be analogously understood as the strong coupling limit of the auxiliary topological string theory. Consequently, we believe that a proper identification with the Kaluza-Klein production in a dual gravity theory (see Gabriel_2000 ; Friedmann:2002gx ; Russo:2009ga for earlier works) may offer key insights into the subject.

### v2

@@ 3.4.2 Alternative computation of the Schwinger determinant
Subsequently, we can add to the integration contour the semi-circle at infinity in the upper half plane since it does not contribute to the complex integral. [^21] This is true, in general, only if $\text{Re}\,\alpha\neq 0$. See Section 4.3.1 for details on this point. Finally, by gluing the integrals in (3.46) to avoid the pole at the origin, we construct a closed path in the complex $s$-plane (see Figure 5), thereby enabling us to rewrite (3.26) as follows [^22] We emphasize that a similar approach to obtaining the full contribution of a given D0-D2 bound state to the generalized prepotential was recently proposed in Hattab:2024ewk .
Figure 5: Integral contour in the complex $s$-plane employed to evaluate the integral (3.47). The singularities located along the imaginary axis are associated to non-perturbative D0-brane effects, whereas the real poles correspond to the perturbative piece (3.28).
@@
which can be finally evaluated upon using the residue theorem. Interestingly, there are two kinds of poles that contribute to the integral (3.47). On the one hand, those occurring along the real axis $s=k\in\mathbb{Z}$ provide for the perturbative piece already computed in (3.28). On the other hand, the poles located at $s=\frac{2\pi in}{\alpha}$ account for non-perturbative D0-brane corrections. Hence, as a final result one obtains
@@
Let us take the opportunity here to stress that the fact that both prescriptions to compute the perturbative and non-perturbative quantum contributions to the generalized prepotential due to the D0-brane tower agree rests, at the end of the day, on us being able to deform the contours for positive/negatively charged states, as well as to add the arc at infinity shown in Figure 5 without any additional cost. Importantly, though, this may not always be necessarily the case, which ultimately depends on the complex phase that the expansion parameter $\alpha$ exhibits (cf. footnote 21). We will elaborate further on this topic later on in Section 4.3.1.
@@ 4 The Fate of Other BPS Black Hole Systems
To accomplish this, we analyze in Section 4.1 another BPS system involving D2- and D6-brane charge. The reason for selecting this family of solutions will become clear along the way. Therefore, following the same strategy as in the previous chapter, we first describe these black holes from the perspective of the two-derivative supergravity theory. Subsequently, in Section 4.2, we repeat the analysis taking into account the effect of the higher-derivative F-terms introduced around eq. (2.5). A key difference between this configuration and the D0-D2-D4 black hole system is that the expansion parameter controlling the quantum deformations of the theory is now purely imaginary. However, as we argue in Section 4.2.2, by appropriately choosing the gauge charges one is able to probe the pathological regime $|\alpha|=\mathcal{O}(1)$ here as well. Nevertheless, in Section 4.3, we show that it is still possible to include highly non-local effects due to the tower of D0-branes which allow us to resum their Schwinger contribution in an exact (i.e., non asymptotic) way. Crucially, we also find that non-perturbative effects are absent in this class of backgrounds, and hence do not modify the attractor solutions nor the entropy. This nicely matches with the observations made in previous sections.
@@
Before proceeding with our discussion, let us briefly summarize here our findings. First of all, we observe that, as soon as we turn on the D6-brane charge in the system, the solution can no longer explore the genuine 5d regime, which we recall corresponds to the $|\alpha|\to\infty$ limit. This can be readily checked from the attractor mechanism, even at two-derivative level. The latter imposes an upper bound on the stabilized value for the (absolute value of) expansion parameter $\alpha$, whenever $p^{0}\neq 0$. Alternatively, from the M-theory perspective, one can argue that having non-trivial D6-brane charge is equivalent to introducing a background Taub-NUT geometry whose center coincides with that of the black hole (see Appendix B for details). This implies, among other things, that in the attractor locus at most $r_{h}\sim r_{5}$} can occur, thus preventing us from performing a clean matching with the putative 5d index, similarly to what we did in Section 3.3.2. Still, by analyzing a representative example introduced in Section 4.1, we confirm that an EFT transition must happen when $|\alpha|$ becomes of order one, which is signaled by an apparent singular behavior exhibited by the asymptotic series of corrections to, e.g., the entropy. This problem can be cured by resorting to the uplifted 5d theory, thereby including highly non-local effects involving the extra circular direction. In any event, very remarkably, we find that non-perturbative contributions in the Scwhinger integral also decouple from this kind of solutions, as it was the case in the simpler D0-D2-D4 system. Finally, in Section 4.3.1 we point out that a simple alternative approach to compute the non-perturbative pair-production-like effects based on Cauchy (cf. Section 3.4.2) seems difficult to apply herein, thus requiring from a special treatment.
@@ 4.1.1 The two-derivative analysis
The family of black hole solutions with $p^{0}=0$ introduced in Section 3.1 is rather special, since they exhibit an explicit attractor for any combination of the remaining quantized charges (see, however, footnote 12). On the other hand, in the most general situation with non-trivial $p^{0},q_{0}$ and arbitrary $(q_{a},p^{b})$-charges, the system is characterized instead by a set of algebraic quadratic equations which may or may not have a real solution, even at the classical level of approximation Shmakova:1996nz . Therefore, in order to provide yet another instance where quantum corrections to the entropy $\mathcal{S}_{\rm BH}$ can be determined and subsequently studied, we consider in what follows the restricted case of 4d black holes with $p^{a}=0$, i.e., with no D4-brane charge. The reason for this choice is twofold. First, such solutions —which are shown to exist for any relative value of the horizon and asymptotic M-theory radii Gauntlett:2002nw ; Gaiotto:2005gf — can be more easily analyzed than their most general counterparts. Second, they uplift to five-dimensional spinning BPS black holes at the center of a Taub-NUT geometry (cf. Appendix B), in contrast to the D0-D2-D4 configuration, which instead corresponds to 5d black strings extended along the compactification circle.
@@
a condition that can be easily attained upon imposing the hierarchy $q_{a}\gg p^{0}$. [^24] Notice that, given a solution $\{x^{a}\}$ of (4.4), one may obtain similar ones upon rescaling $q_{a}\to k^{\beta}q_{a}$ and $p^{0}\to k^{\gamma}p^{0}$}, which results into $x^{a}\to k^{\frac{\beta-\gamma}{2}}x^{a}$. Hence, by taking $\beta>\gamma$ and $k\gg 1$, one can easily achieve (4.10). Notice that, similarly to what happened with the D0-D2-D4 system (cf. (3.9)), at this level of approximation we do not need to specify the asymptotics of $CX^{0}$, which in the present case is entirely determined by the D6-brane charge $p^{0}$. The latter turns out to control the importance of the relevant, perturbative quantum corrections, as we discuss next.
@@ 4.2.1 Including higher-derivative corrections
$|\mathscr{Z}|^{2}$ $=-\frac{4}{3}Y^{a}\left(q_{a}-\frac{1}{12p^{0}}c_{2,a}\right)+p^{0}\,\text{Re}\,G_{0}\,$, (4.13)
@@
$\mathcal{S}_{\rm BH}=-\frac{4}{3}\pi Y^{a}\left(q_{a}+\frac{1}{6p^{0}}c_{2,a}\right)+\frac{\chi_{E}(X_{3})(p^{0})^{2}}{4(2\pi)^{2}}\sum_{g=0,2,3,\ldots}(-1)^{g}c^{3}_{g-1}\,|\alpha|^{2g}\,.$ (4.14)
@@ 4.2.2 The transition regime
Let us also remark that the absence of an asymptotic five-dimensional regime prevents the existence of a genuine 5d configuration in the decompactified theory that could be used to ‘glue’ with the D2-D6 black hole across the transition regime. This observation was already clear from the perspective of the classical, two-derivative theory, given that the KK monopoles sourcing the Taub-NUT charge are topological solitons that exist only in presence of compact directions, see Appendix B.1. Furthermore, we can confirm now that the previous conclusion is not modified upon including the relevant set of quantum corrections (perturbative and non-perturbative). Therefore, we cannot test our quantum-corrected entropy in the same way as in Section 3 for the D0-D2-D4 black hole, where we ultimately recovered the (regularized) entropy of the 5d black string carrying M2, M5 and KK charges by taking the $\alpha\to\infty$ limit.
@@ 4.3 Non-local and non-perturbative effects
Figure 7: Numerical plot of $\mathcal{I}(|\alpha|)$ as a function of the (purely imaginary) expansion parameter $|\alpha|$. The actual computation corresponds to the ‘renormalized’ expression displayed in (4.22). Notice that the integral is convergent and finite for every $|\alpha|\in[0,2)$ within its physical domain (see footnote 27).
where we have explicitly separated the contributions for $g\leq 1$ and $g>1$, subsequently introduced the integration parameter $s=\pi n|\alpha|\tau$ and finally performed the summation of the geometric series in $e^{-4\pi s/|\alpha|}$. The terms subtracted in the parenthesis correspond to the regularization of the pole at the origin, which allow us to safely remove the cutoff $0^{+}$ in (4.21) (see, for instance, Schwartz:2014sze ). We note, in particular, that the integral above can be easily checked to be convergent (cf. Figure 7) and moreover defines the exact, resummed version of the asymptotic series (4.12a). To see this, one may insert back the Laurent series for $\text{csch}^{2}(x)$ at $x=0$ (cf. eq. (3.27)) in (4.22), subsequently exchange the order of summation and integration, and finally perform the integral for each term independently. For completeness, we show in Figure 8 a numerical evaluation [^26] The authors would like to thank Alessandro Lenoci for useful explanations on how to perform the high-precision numerical evaluation. of the error made by the asymptotic approximation to the one-loop determinant when $|\alpha|\ll 1$. [^27] As a side note, let us point out that the integral (4.22) is well-behaved and monotonic for $|\alpha|>2$ as well. It becomes negative around $|\alpha|\simeq 8$, and moreover behaves like $|\alpha|^{3}$ for $|\alpha|\gg 1$, similarly to what happened with the non-perturbative contribution $\mathcal{I}^{(np)}(\alpha)$ in the D0-D2-D4 case, cf. eq. (3.45). Let us also mention that the absence of additional non-perturbative terms in $G(Y^{0},\Upsilon)$ can be understood as well from the fact that the series (4.12) are, in this case, alternating and thus Borel summable (see Appendix A for details), contrary to what happened in the wrapped black string background, cf. eq. (3.16).
@@ 4.3.1 Challenges and Obstructions in the Cauchy formulation
$\mathcal{I}^{(np)}(\alpha)\,$ $\sim\,-2\pi i\alpha\sum_{n=1}^{\infty}\frac{1}{n}e^{-\frac{4n\pi^{2}}{\alpha}}\,$, (4.24b)
@@
Figure 9: Integral contour in the complex $s$-plane employed to evaluate the one-loop integral (4.23). The non-perturbative singularities no longer lie along the imaginary axes if $\alpha$ has a non-vanishing complex phase $\theta_{\alpha}$. The real poles still correspond to the perturbative piece. In the limit of $\text{Re}\,\alpha\to 0$, all the poles therefore become real. However, for the rational values of $\alpha$ enforced by the attractor mechanism, perturbative and non-perturbative singularities do not coincide.
The divergences we are encountering in the case of $\text{Re}(\alpha)=0$ signal that we should not be allowed to use anymore the residue theorem when evaluating (4.23). Indeed, in its standard formulation (see, e.g., Ahlfors1966 ), one considers a finite arc as well as finitely many poles. One can then formally extend the computation to the case of infinitely many isolated singularities by considering a discrete family of contours $\{\mathcal{C}_{N}\}$ which enclose $N$ such poles, and subsequently take the limit $N\to\infty$. This process is well-defined if and only if the limit exists, i.e., provided the series of residues converges. For $\alpha$ imaginary this does not happen, as we just discussed, and thus we cannot use the residue theorem to evaluate the integral. Notice that this oscillatory behavior can be ultimately traced back to the contributions associated to the infinitesimal semi-circles surrounding each pole along the real axis. In essence, what happens is that, since both sets of singularities are dense with respect to each other, the small arcs around the ‘quasi-poles’ get an enhancement —due to the closeness to the nearest singularity— that grows as we go towards real-infinity, as per (4.26).
In addition, it is worth highlighting that, from this perspective, one can also understand the origin of the harmonic-like behavior exhibited by the series of residues (4.25). Recall that the starting point was eq. (3.26), which does require special care when $\alpha$ is purely imaginary. Whenever this happens, what one can do is to consider a shift so as to move the non-perturbative singularities infinitesimally away from the real axis, as we did in (3.46). While this deformation may seem sufficient to eliminate the divergences caused by going exactly through poles, it is not enough to fully regularize the expression. To see this, let us consider the integrand of (3.46a) with $\alpha=i|\alpha|$, and where the non-perturbative poles properly are shifted as explained before, namely
@@
where we substituted $\alpha=i|\alpha|$ and we defined $\tau=is$, cf. footnote 15. Note that the above expression is clearly reminiscent of the exact result derived in eq. (4.22), and it is moreover well-defined and convergent (except for the pole at $\tau=0$, which must be carefully dealt with as in the rest of this work). Furthermore, from this point of view it becomes clear that the pathologies we referred to in our previous discussion would be absent in (4.30) and hence only appear if we close the contour by adding the arc at infinity, whereby picking up the residues displayed in (3.48).
@@ 5 Conclusions and Outlook
In this work, we have revisited the role played by macroscopic quantum corrections to the black hole entropy in 4d $\mathcal{N}=2$ supersymmetric effective field theories arising from Type IIA string theory compactified on Calabi–Yau threefolds. More precisely, by considering an infinite series of higher-derivative F-terms, we examined the most relevant perturbative and non-perturbative contributions to the supersymmetric black hole index close to the large volume point, emphasizing the interplay between different limiting (dual) descriptions in four and five spacetime dimensions. In particular, we obtained an explicit and well-defined, analytic expression for the entropy which allows us to track the underlying physical system across the transition between the two different field-theoretic descriptions. As a byproduct, this study highlights the importance of EFT transitions in understanding quantum-gravitational aspects of black hole thermodynamics.
@@
One of the main conclusions of our investigation is that the aforementioned quantum corrections to the entropy exhibit distinct behaviors as the characteristic size of the black hole approaches the Kaluza-Klein scale, where the effects of the extra dimensions become relevant. In particular, we identified a ‘transition regime’ (cf. Section 3.2.2) where an infinite number of local higher-curvature and higher-derivative operators seem to induce pathological contributions to the macroscopic black hole entropy, signaling the failure of the 4d $\mathcal{N}=2$ EFT to correctly describe such configurations. Indeed, the corrections organize into an asymptotic series with expansion parameter $\alpha$ which can be interpreted as the ratio of the inverse D0-brane mass to the black hole radius. We showed that one can regularize the series trough a resummation procedure and how the highly non-local perturbative effects induced by the infinite tower of Kaluza-Klein states (identified herein as D0 bound states) plays a crucial role in regulating certain ultra-violet divergences naively exhibited by the entropy function. This was explicitly analyzed for two BPS configurations, namely the D0-D2-D4 (Section 3) and the D2-D6 (Section 4) black holes. Furthermore, in a sense, these two systems are somewhat complementary. From a more technical point of view, the resummation procedure that was used for the former black hole can be readily extended for more general charge configurations. Crucially, though, it fails precisely when we turn off D0- and D4-brane charges, which is equivalent to the regime of purely imaginary $\alpha$, see discussion in Section 4.3.1. Therefore, in Section 4, we not only studied a second example but also clarified how to extend the resummation prescription employed in Section 3 to this case as well. Moreover, as discussed in Section 3.3.2, by lifting the solution to five dimensions and enforcing a parametrically small horizon radius compared to the size of the internal circle (whenever possible), one finds perfect agreement with the microscopic counting Maldacena:1997de ; Vafa:1997gr ; Harvey:1998bx and one-loop exact computations in the dual M-theory description Castro:2007sd . The transition regime can be crossed completely by taking some appropriate limit only for the D0-D2-D4 system. Indeed, the aforementioned black hole becomes a 5d black string wrapped along the extra compact direction, which grows indefinitely in the decompactification limit. Instead, the D2-D6 solutions uplifts to a five-dimensional black hole with Kaluza–Klein monopole charge, and it exists as long as the extra dimension is strictly compact. In this latter case, we can explore the transition regime; however, a topological obstruction prevents us from taking the full decompactification limit for this class of configurations. Remarkably, we also found that additional non-perturbative corrections —which are oftentimes present and may be related to non-trivial pair production rate of Kaluza-Klein gravitons (and superpartners thereof) in the anti-self-dual graviphoton constant background close to the black hole horizon— seem to not modify the most relevant physical properties of the solutions considered in this work, in particular their associated entropy (cf. Sections 3.4 and 4.3). On the one hand, this further supports the consistency of the analysis carried out here. On the other hand, it raises the question of whether the same phenomenon might also occur for more general BPS black holes that can be constructed in the 4d theory. A more detailed investigation of these issues is left for future work Castellano:2025yur .
@@
On another note, one may hope to be able to obtain from this perspective further insights into the non-perturbative behavior of certain topological string theories, which are known to capture the same prepotential controlling the higher-derivative corrections to the entropy Ooguri:2004zv . In fact, we saw that in the D0-D2-D4 black hole background, namely the one associated to the unique system that is able to explore the genuine 5d regime, there should be a priori certain non-perturbative contributions to the generalized prepotential. These can be equivalently determined via a careful study of a one-loop integral associated to BPS states in M-theory Gopakumar:1998ii ; Gopakumar:1998jq ; Pasquetti:2010bps (see also Hattab:2024ewk ; Hattab:2024ssg for recent related results). Hence, even though the latter did not ultimately affect any of the black hole observables we actually cared about in this work, they certainly exhibited interesting behaviors, particularly so along the decompactification limit. Interestingly, this regime can be analogously understood as the strong coupling limit of the auxiliary topological string theory. Consequently, we believe that a proper identification with the Kaluza-Klein production in a dual gravity theory (see Gabriel_2000 ; Friedmann:2002gx ; Russo:2009ga for earlier works) may offer key insights into the subject.